Exercise 15-10 For a recent 2-year period, the balance sheet of Martinez Company showed the following stockholders equity data at December 31 (in millions) 2017 2016 Additional paid-in capital Common stock Retained earnings Treasury stock $910 675 7,210 1,833 $6,962 $846 663 5,300 Total stockholders equity $5,810 Common stock shares issued Common stock shares authorized Treasury stock shares 225 500 39 221 500 27 (a) Answer the following questions. (1) What is the par value of the common stock? (Round par value to 2 decimal places, e.g. $3.15.) Par value of common stock (2) What is the cost per share of treasury stock at December 31, 2017, and at December 31, 2016? December 31, 2017 December 31, 2016 Cost per share of Treasury stock
Answer #1

Par value of common stock = 675/225 = 3 per share

Cash per share of treasury stock 2016 = 999/27 = 37

Cash per share of treasury stock 2017 = 1833/39 = 47

Prepare shareholder's equity section :

Contributed capital
Paid in capital
Common stock 675
Paid in capital in excess of par value 910
Total paid in capital 1585
Retained earnings 7210
Total 8795
Less: Treasury stock (1833)
Total shareholder's equity 6962